A Business Firms sells a good at the price of Rs 450. The firm has decided to reduce the price of good to Rs 350. Consequently , the quantity demanded for the good rose from 25,000 units to 35,000 units . Calculate the price elasticity of demand
1
Expert's answer
2020-11-25T11:51:00-0500

P1 = 450

P2 = 350

Q1 = 25000

Q2 = 35000

Percentage change in quantity "= \frac{Q_2 – Q_1}{\frac{Q_2 + Q_1}{2}} \times 100"

Percentage change in quantity "= \frac{35000 – 25000}{\frac{35000 + 25000}{2}} \times 100 = \frac{100}{3}"

Percentage change in price "= \frac{P_2 – P_1}{\frac{P_2 + P_1}{2}} \times 100"

Percentage change in price "= \frac{350 – 450}{\frac{350 + 450}{2}} \times 100 = \frac{100}{4}"

Price elasticity of demand "= \frac{Percentage \;change \;in\; quantity}{Percentage\; change\; in\; price}"

Price elasticity of demand "= \frac{100}{3} \times \frac{4}{100} = \frac{4}{3} = 1.333"

The price elasticity of demand is 1.333.